# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 7.3 Stabilizers of rank $1$ vertices are virtually cyclic
Theorem 7.10 does not hold if $L$ is not assumed to have bounded torsion. The simplest example is to take $L=\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$. Then $L$ and $\mathbb{Z}$ have orbit equivalent actions on $X=\{0,1\}^{\mathbb{N}}$ (equipped with the product measure of the uniform measure on $\{0,1\}$). Here $\mathbb{Z}$ acts as the odometer, and $L$ acts by coordinatewise addition; two elements are in the same orbit if and only if they have the same tail. Letting $c:L\times X\to\mathbb{Z}$ be the orbit equivalence cocycle, it is easy to see that $|c(s,\cdot)|$ is bounded for every $s\in L$.
@@ 8 Controlling the factor actions
Throughout the section, we let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<\infty$, and $H$ be a countable group with bounded torsion, such that there exists an $(L^{1},L^{0})$-measure equivalence coupling $\Omega$ from $H$ to $G$. Let $\hat{\Omega}$ be the measure equivalence coupling between $H$ and $\hat{G}$ defined by letting $\hat{\Omega}=(\hat{G}\times\Omega)/G$ as in Section 7.1. Let $\iota:H\to\mathrm{Aut}(\Gamma^{e})$ and $\theta:\hat{\Omega}\to\mathrm{Aut}(\Gamma^{e})$ be the maps given by Lemma 6.6 through the canonical isomorphism $\mathrm{Aut}(\mathbb{B})\simeq\mathrm{Aut}(\Gamma^{e})$ recalled in Section 2.2. In particular $H$ acts on $\Gamma^{e}$, and therefore $H$ acts on $G$ by flat-preserving bijections, see Section 2.3.
@@
The goal of this section is to understand the action $H_{\mathsf{v}}\curvearrowright P_{\mathsf{v}}$ to the extent that we can connect with the assumptions of Theorem 3.2.
@@ 8.1 Finite generation of $H_{\mathsf{v}}$
The main goal of this subsection is to prove that $H_{\mathsf{v}}$ is finitely generated (Lemma 8.4 below).
@@
Recall that the $H$-action on $\mathbb{B}$ can be equivalently viewed as an $H$-action on $G$ by flat-preserving bijections. Using this viewpoint, for every vertex $v\in V\mathbb{B}$ with associated standard flat $F$, the stabilizer $H_{v}$ coincides with the stabilizer $H_{F}$ of $F$ for this $H$-action on $G$. We will use this viewpoint throughout the proof. We first claim that for any standard flat $F^{\prime}\subset F$, the group $H_{F^{\prime}}$ has a finite index subgroup $H^{0}_{F^{\prime}}$ which is contained in $H_{F}$, and acts on $F^{\prime}$ with finitely many orbits. Indeed, let $v_{F^{\prime}}$ be the vertex of $\mathbb{B}$ associated to $F^{\prime}$. Lemma 6.8 ensures that $H_{F^{\prime}}$ acts on the set of rank 0 vertices of $\mathbb{B}$ that are smaller than $v_{F^{\prime}}$ with finitely many orbits. Therefore $H_{F^{\prime}}$ acts on $F^{\prime}$ with finitely many orbits. Any element in $H_{F^{\prime}}$ sends $F$ to another standard flat containing $F^{\prime}$. As there are only finitely many standard flats containing $F^{\prime}$, we can find a finite index subgroup $H^{0}_{F^{\prime}}\subseteq H_{F^{\prime}}$ that stabilizes $F$. The action $H^{0}_{F^{\prime}}\curvearrowright F^{\prime}$ still has finitely many orbits, thus the claim is proved. We now prove the lemma by induction on the rank of vertices of $\mathbb{B}$. Stabilizers in $H$ of rank 00 vertices are finite (Lemma 6.8), and the case of rank 1 vertices is given by Proposition 7.1. Let now $v\in V\mathbb{B}$ be a vertex of rank $n\geq 2$, corresponding to a standard flat $F$, and assume by induction that the lemma is proven for all vertices of rank at most $n-1$. Take a standard line $\ell\subset F$. By the above claim, there is a finite index subgroup $H^{0}_{\ell}\subseteq H_{\ell}$ that stabilizes both $F$ and $\ell$, and whose action on $\ell$ has finitely many orbits. Let $\mathcal{C}$ be the set of all standard flats in $F$ of dimension $\dim F-1$ that intersect $\ell$ in exactly one point. As the $H$-action on $F$ is flat-preserving, the group $H_{\ell}^{0}$ permutes elements in $\mathcal{C}$, and the permutation action $H_{\ell}^{0}\curvearrowright\mathcal{C}$ has finitely many orbits. Let $\{F_{1},\dots,F_{k}\}$ be a set consisting of exactly one representative in each orbit of this permutation action. By the above claim, for each $i\in\{1,\dots,k\}$, there is a finite index subgroup $H^{0}_{F_{i}}\subseteq H_{F_{i}}$ that stabilizes both $F$ and $F_{i}$, and acts on $F_{i}$ with finitely many orbits. Let $K_{i}$ be a finite subset of $F_{i}$ such that $H^{0}_{F_{i}}K_{i}=F_{i}$. Let $K=\cup_{i=1}^{k}K_{i}$ and let $H^{\prime}$ be the subgroup of $H_{F}$ generated by $H^{0}_{\ell}$ and $\{H^{0}_{F_{i}}\}_{i=1}^{k}$. As $H^{0}_{\ell}$ and $H^{0}_{F_{i}}$ are finitely generated by induction, it follows that $H^{\prime}$ is finitely generated. Moreover, our construction implies that $H^{\prime}K=F$. Thus for any $h\in H_{F}$, there exists $h^{\prime}\in H^{\prime}$ such that $h^{\prime}hK\cap K\neq\emptyset$. On the other hand, as the action $H_{F}\curvearrowright F$ has finite stabilizers (Lemma 6.8) and $K$ is finite, there are only finitely many elements $h\in H_{F}$ such that $hK\cap K\neq\emptyset$. Thus $H^{\prime}$ has finite index in $H_{F}$, and hence $H_{F}$ is finitely generated. ∎
@@
If $F\subseteq P_{\mathsf{v}}$ is a standard flat (corresponding to a vertex of $\mathbb{B}(\mathsf{v})$), then $F$ is contained in only finitely many regions of the form $P_{\mathsf{w}}$ with $\mathsf{w}\in V\Gamma^{e}$. Then $\mathrm{Stab}_{H}(F)$ has a finite index subgroup preserving $P_{\mathsf{v}}$. By the claim in the previous paragraph, it follows that $\mathrm{Stab}_{H_{\mathsf{v}}}(F)$ is a finite-index subgroup of $\mathrm{Stab}_{H}(F)$, and is thus finitely generated by Lemma 8.2. More generally, by the same argument (and using Corollary 8.3), for any cube $C$ in $\mathbb{B}(\mathsf{v})$, the $H_{\mathsf{v}}$-stabilizer of $C$ has finite index in its $H$-stabilizer, hence is finitely generated. Now $H_{\mathsf{v}}$ acts by cubical automorphisms on the simply connected complex $\mathbb{B}(\mathsf{v})$ cocompactly (Lemma 8.1(3)), with finitely generated cell stabilizers, so it is finitely generated by [Bro84, Theorem 1]. ∎
@@ 8.2 Commensurability of stabilizers
For Fact 1, note that the action $G_{\mathsf{v}}\curvearrowright P_{\mathsf{v}}$ is transitive and respects the splitting $P_{\mathsf{v}}\cong Z_{\mathsf{v}}\times\mathcal{L}_{\mathsf{v}}$, so the action $G_{\mathsf{v}}\curvearrowright Z_{i}$ is transitive for $i=1,2$. Hence the same holds for the bigger group $\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$. Fact 2 follows from the discussion of stabilizers of factor actions before the proposition. As before, let $\hat{\Omega}$ be a measure equivalence coupling between $\hat{G}$ and $H$, and let $\theta:\hat{\Omega}\to\mathrm{Aut}(\mathbb{B})\cong\mathrm{Aut}(\Gamma^{e})$ be a measurable $(\hat{G}\times H)$-equivariant map. As recalled in the introductory paragraph of Section 8, the space $\hat{\Omega}_{\mathsf{v}}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v}}(\mathbb{B}))$ is a measure equivalence coupling between $\hat{G}_{\mathsf{v}}$ and $H_{\mathsf{v}}$. Hence $\hat{\Omega}_{\mathsf{v}}$ is also a measure equivalence coupling between $G_{\mathsf{v}}$ and $H_{\mathsf{v}}$. Let $\hat{\Omega}_{\mathsf{v},x}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B}))$. Since the orbits of $x$ under $\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$ and $G_{\mathsf{v}}$ coincide (Fact 1 above), it follows from Corollary 4.4 (applied with $K=Z_{i}$, with $L=\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$, with $\mathsf{G}=G_{\mathsf{v}}$, and with $\hat{\Omega}_{\mathsf{v}}$ in place of $\Omega$) that $\hat{\Omega}_{\mathsf{v},x}$ is a measure equivalence coupling between $G_{\mathsf{v},x}$ and $H_{\mathsf{v},x}$. We want to apply Lemma 4.9 with $\mathsf{G}=G_{\mathsf{v},x},\mathsf{H}=H_{\mathsf{v},x},\Sigma=\hat{\Omega}_{\mathsf{v},x}$ and $\mathsf{H}^{\prime}=H_{\mathsf{v},x,y}$. It remains to find $\Sigma^{\prime}$ with desired properties. For the following discussion, we refer to the statement of Lemma 6.6 for our convention of how $\mathsf{G}$ and $\mathsf{H}$ act on $\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B})$. Given $z\in Z_{i}$, we let $\mathrm{Aut}_{\mathsf{v},x,y\to z}(\mathbb{B})$ be the Borel subset of $\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B})$ consisting of all automorphisms that send $y$ to $z$, and let $\hat{\Omega}_{\mathsf{v},x,y\to z}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v},x,y\to z}(\mathbb{B}))$. Then $\hat{\Omega}_{\mathsf{v},x}=\sqcup_{z\in Z_{i}}\hat{\Omega}_{\mathsf{v},x,y\to z}$. Therefore, we can (and will) choose $z\in Z_{i}$ such that $\mu(\hat{\Omega}_{\mathsf{v},x,y\to z})>0$. We take $\Sigma^{\prime}=\hat{\Omega}_{\mathsf{v},x,y\to z}$. It is invariant under $G_{\mathsf{v},x,z}$ and $\mathsf{H}^{\prime}=H_{\mathsf{v},x,y}$, and the former group is equal to $\mathsf{G}=G_{\mathsf{v},x}$ by Fact 2. Now take $h\in\mathsf{H}\setminus\mathsf{H}^{\prime}$. Then $h(x)=x$ and $h(y)\neq y$. Then $h\Sigma^{\prime}=\hat{\Omega}_{\mathsf{v},x,h(y)\to z}$, which is disjoint from $\hat{\Omega}_{\mathsf{v},x,y\to z}$, as desired. ∎
@@ 8.3 From commensurated to normal
Let $\mathsf{v}\in V\Gamma^{e}$. Let $\ell_{1},\ell_{2}$ be two $\mathsf{v}$-lines in the same $H_{\mathsf{v}}$-orbit. Let $A$ be any finite-index infinite cyclic subgroup of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\cap\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$. Then $[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1}):A]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2}):A]$.
@@
Recall that every $\mathsf{v}$-line $\ell$ determines a rank $1$ vertex $v_{\ell}$ in $\mathbb{B}$, and vertices of $\ell$ correspond to rank 00 vertices adjacent to $v_{\ell}$. Therefore, for every $\mathsf{v}$-line $\ell$, the action of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ on $\ell$ has finite stabilizers (Lemma 6.8) and finitely many orbits (Lemma 8.1). As $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})$ and $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$ are commensurable (Proposition 8.5 applied to $Z_{2}=\mathcal{L}_{\mathsf{v}}$), and $A$ is torsion-free and has finite index in $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\cap\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$, it follows that the actions $A\curvearrowright\ell_{1}$ and $A\curvearrowright\ell_{2}$ are free with finitely many orbits. As $A\subseteq H_{\mathsf{v}}$, the $A$-action on $P_{\mathsf{v}}=Z_{\mathsf{v}}\times\mathcal{L}_{\mathsf{v}}$ preserves the product structure. Therefore the two actions $A\curvearrowright\ell_{1}$ and $A\curvearrowright\ell_{2}$ have the same number of orbits. Let $h\in H_{\mathsf{v}}$ be such that $h(\ell_{1})=\ell_{2}$. Then $h$ induces a bijection $h_{\ast}:\mathcal{O}_{1}^{j}\mapsto\mathcal{O}_{2}^{j}$ between the set of orbits $\{\mathcal{O}^{1}_{1},\dots,\mathcal{O}^{k}_{1}\}$ of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\curvearrowright\ell_{1}$, and the set of orbits $\{\mathcal{O}^{1}_{2},\dots,\mathcal{O}^{k}_{2}\}$ of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})\curvearrowright\ell_{2}$. In addition $h_{\ast}$ preserves the (finite) cardinalities of stabilizers, i.e. the $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})$-stabilizer of any point in $\mathcal{O}_{1}^{j}$ has the same cardinality (denoted by $k_{j}$) as the $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$-stabilizer of any point in $\mathcal{O}_{2}^{j}$. Now, for every $i\in\{1,2\}$, the number of $A$-orbits on $\ell_{i}$ is equal to $[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{i}):A]\sum_{j=1}^{k}\frac{1}{k_{j}}$. Therefore $[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1}):A]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2}):A]$. ∎

### v2

@@ 7.3 Stabilizers of rank $1$ vertices are virtually cyclic
Theorem 7.11 does not hold if $L$ is not assumed to have bounded finite subgroups. The simplest example is to take $L=\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$. Then $L$ and $\mathbb{Z}$ have orbit equivalent actions on $X=\{0,1\}^{\mathbb{N}}$ (equipped with the product measure of the uniform measure on $\{0,1\}$). Here $\mathbb{Z}$ acts as the odometer, and $L$ acts by coordinatewise addition; two elements are in the same orbit if and only if they have the same tail. Letting $c:L\times X\to\mathbb{Z}$ be the orbit equivalence cocycle, it is easy to see that $|c(s,\cdot)|$ is bounded for every $s\in L$.
@@ 8 Controlling the factor actions
Throughout the section, we let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<\infty$, and $H$ be a countable group with bounded finite subgroups, such that there exists an $(L^{1},L^{0})$-measure equivalence coupling $\Omega$ from $H$ to $G$. Let $\hat{\Omega}$ be the measure equivalence coupling between $H$ and $\hat{G}$ defined by letting $\hat{\Omega}=(\hat{G}\times\Omega)/G$ as in Section 7.1. Let $\iota:H\to\mathrm{Aut}(\Gamma^{e})$ and $\theta:\hat{\Omega}\to\mathrm{Aut}(\Gamma^{e})$ be the maps given by Lemma 6.6 through the canonical isomorphism $\mathrm{Aut}(\mathbb{B})\simeq\mathrm{Aut}(\Gamma^{e})$ recalled in Section 2.2. In particular $H$ acts on $\Gamma^{e}$, and therefore $H$ acts on $G$ by flat-preserving bijections, see Section 2.3.
@@
The goal of this section is to understand the action $H_{\mathsf{v}}\curvearrowright P_{\mathsf{v}}$ to the extent that we can connect with the assumptions of Theorem 3.2. We will first prove a few group theoretical properties of $H_{\mathsf{v}}$, namely $H_{\mathsf{v}}$ is finitely generated (Section 8.1), and $H_{\mathsf{v}}$ contains a commensurated $\mathbb{Z}$ subgroup (Section 8.2), which is actually virtually normal (Section 8.3), before we prove the main lemma (Lemma 8.9) on the action $H_{\mathsf{v}}\curvearrowright P_{\mathsf{v}}$ in Section 8.4.
@@ 8.1 Finite generation of $H_{\mathsf{v}}$
The main goal of this subsection is to prove that $H_{\mathsf{v}}$ is finitely generated (Lemma 8.4 below). Recall that finite generation has already been established for $H$-stabilizers of vertices of rank at most 1 in $\mathbb{B}$ (Proposition 7.1), and this will be used to prove the finite generation of $H_{\mathsf{v}}$.
@@
Recall that the $H$-action on $\mathbb{B}$ can be equivalently viewed as an $H$-action on $G$ by flat-preserving bijections. Using this viewpoint, for every vertex $v\in V\mathbb{B}$ with associated standard flat $F$, the stabilizer $H_{v}$ coincides with the stabilizer $H_{F}$ of $F$ for this $H$-action on $G$. We will use this viewpoint throughout the proof.
@@
We first claim that for any standard flat $F^{\prime}\subset F$, the group $H_{F^{\prime}}$ has a finite index subgroup $H^{0}_{F^{\prime}}$ which is contained in $H_{F}$, and acts on $F^{\prime}$ with finitely many orbits. Indeed, let $v_{F^{\prime}}$ be the vertex of $\mathbb{B}$ associated to $F^{\prime}$. Lemma 6.8 ensures that $H_{F^{\prime}}$ acts on the set of rank 0 vertices of $\mathbb{B}$ that are smaller than $v_{F^{\prime}}$ with finitely many orbits. Therefore $H_{F^{\prime}}$ acts on $F^{\prime}$ with finitely many orbits. Any element in $H_{F^{\prime}}$ sends $F$ to another standard flat containing $F^{\prime}$. As there are only finitely many standard flats containing $F^{\prime}$, we can find a finite index subgroup $H^{0}_{F^{\prime}}\subseteq H_{F^{\prime}}$ that stabilizes $F$. The action $H^{0}_{F^{\prime}}\curvearrowright F^{\prime}$ still has finitely many orbits, thus the claim is proved.
We now prove the lemma by induction on the rank of vertices of $\mathbb{B}$. Stabilizers in $H$ of rank 0 vertices are finite (Lemma 6.8), and the case of rank 1 vertices is given by Proposition 7.1. Let now $v\in V\mathbb{B}$ be a vertex of rank $n\geq 2$, corresponding to a standard flat $F$, and assume by induction that the lemma is proven for all vertices of rank at most n1n-1. Take a standard line $\ell\subset F$. By the above claim, there is a finite index subgroup $H^{0}_{\ell}\subseteq H_{\ell}$ that stabilizes both $F$ and $\ell$, and whose action on $\ell$ has finitely many orbits. Let $\mathcal{C}$ be the set of all standard flats in $F$ of dimension $\dim F-1$ that intersect $\ell$ in exactly one point. As the $H$-action on $F$ is flat-preserving, the group $H_{\ell}^{0}$ permutes elements in $\mathcal{C}$, and the permutation action $H_{\ell}^{0}\curvearrowright\mathcal{C}$ has finitely many orbits. Let $\{F_{1},\dots,F_{k}\}$ be a set consisting of exactly one representative in each orbit of this permutation action. By the above claim, for each $i\in\{1,\dots,k\}$, there is a finite index subgroup $H^{0}_{F_{i}}\subseteq H_{F_{i}}$ that stabilizes both $F$ and $F_{i}$, and acts on $F_{i}$ with finitely many orbits. Let $K_{i}$ be a finite subset of $F_{i}$ such that $H^{0}_{F_{i}}K_{i}=F_{i}$. Let $K=\cup_{i=1}^{k}K_{i}$ and let $H^{\prime}$ be the subgroup of $H_{F}$ generated by $H^{0}_{\ell}$ and $\{H^{0}_{F_{i}}\}_{i=1}^{k}$. As $H^{0}_{\ell}$ and $H^{0}_{F_{i}}$ are finitely generated by induction, it follows that $H^{\prime}$ is finitely generated. Moreover, our construction implies that $H^{\prime}K=F$. Thus for any $h\in H_{F}$, there exists $h^{\prime}\in H^{\prime}$ such that $h^{\prime}hK\cap K\neq\emptyset$. On the other hand, as the action $H_{F}\curvearrowright F$ has finite stabilizers (Lemma 6.8) and $K$ is finite, there are only finitely many elements $h\in H_{F}$ such that $hK\cap K\neq\emptyset$. Thus $H^{\prime}$ has finite index in $H_{F}$, and hence $H_{F}$ is finitely generated. ∎
@@
If $F\subseteq P_{\mathsf{v}}$ is a standard flat (corresponding to a vertex of $\mathbb{B}(\mathsf{v})$), then $F$ is contained in only finitely many regions of the form $P_{\mathsf{w}}$ with $\mathsf{w}\in V\Gamma^{e}$. Then $\mathrm{Stab}_{H}(F)$ has a finite index subgroup preserving $P_{\mathsf{v}}$. By the claim in the previous paragraph, it follows that $\mathrm{Stab}_{H_{\mathsf{v}}}(F)$ is a finite-index subgroup of $\mathrm{Stab}_{H}(F)$, and is thus finitely generated by Lemma 8.2. More generally, by the same argument (and using Corollary 8.3), for any cube $C$ in $\mathbb{B}(\mathsf{v})$, the $H_{\mathsf{v}}$-stabilizer of $C$ has finite index in its $H$-stabilizer, hence is finitely generated.
Now $H_{\mathsf{v}}$ acts by cubical automorphisms on the simply connected complex $\mathbb{B}(\mathsf{v})$ cocompactly (Lemma 8.1(3)), with finitely generated cell stabilizers, so it is finitely generated by [Bro84, Theorem 1]. ∎
@@ 8.2 Commensurability of stabilizers
For Fact 1, note that the action $G_{\mathsf{v}}\curvearrowright P_{\mathsf{v}}$ is transitive and respects the splitting $P_{\mathsf{v}}\cong Z_{\mathsf{v}}\times\mathcal{L}_{\mathsf{v}}$, so the action $G_{\mathsf{v}}\curvearrowright Z_{i}$ is transitive for $i=1,2$. Hence the same holds for the bigger group $\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$. Fact 2 follows from the discussion of stabilizers of factor actions before the proposition.
As before, let $\hat{\Omega}$ be a measure equivalence coupling between $\hat{G}$ and $H$, and let $\theta:\hat{\Omega}\to\mathrm{Aut}(\mathbb{B})\cong\mathrm{Aut}(\Gamma^{e})$ be a measurable $(\hat{G}\times H)$-equivariant map. As recalled in the introductory paragraph of Section 8, the space $\hat{\Omega}_{\mathsf{v}}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v}}(\mathbb{B}))$ is a measure equivalence coupling between $\hat{G}_{\mathsf{v}}$ and $H_{\mathsf{v}}$. Hence $\hat{\Omega}_{\mathsf{v}}$ is also a measure equivalence coupling between $G_{\mathsf{v}}$ and $H_{\mathsf{v}}$.
Let $\hat{\Omega}_{\mathsf{v},x}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B}))$. Since the orbits of $x$ under $\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$ and $G_{\mathsf{v}}$ coincide (Fact 1 above), it follows from Corollary 4.5 (applied with $K=Z_{i}$, with $L=\mathrm{Aut}_{\mathsf{v}}(\mathbb{B})$, with $\mathsf{G}=G_{\mathsf{v}}$, and with $\hat{\Omega}_{\mathsf{v}}$ in place of $\Omega$) that $\hat{\Omega}_{\mathsf{v},x}$ is a measure equivalence coupling between $G_{\mathsf{v},x}$ and $H_{\mathsf{v},x}$.
We want to apply Lemma 4.10 with $\mathsf{G}=G_{\mathsf{v},x},\mathsf{H}=H_{\mathsf{v},x},\Sigma=\hat{\Omega}_{\mathsf{v},x}$ and $\mathsf{H}^{\prime}=H_{\mathsf{v},x,y}$. It remains to find $\Sigma^{\prime}$ with desired properties. For the following discussion, we refer to the statement of Lemma 6.6 for our convention of how $\mathsf{G}$ and $\mathsf{H}$ act on $\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B})$.
Given $z\in Z_{i}$, we let $\mathrm{Aut}_{\mathsf{v},x,y\to z}(\mathbb{B})$ be the Borel subset of $\mathrm{Aut}_{\mathsf{v},x}(\mathbb{B})$ consisting of all automorphisms that send $y$ to $z$, and let $\hat{\Omega}_{\mathsf{v},x,y\to z}=\theta^{-1}(\mathrm{Aut}_{\mathsf{v},x,y\to z}(\mathbb{B}))$. Then $\hat{\Omega}_{\mathsf{v},x}=\sqcup_{z\in Z_{i}}\hat{\Omega}_{\mathsf{v},x,y\to z}$. Therefore, we can (and will) choose $z\in Z_{i}$ such that $\mu(\hat{\Omega}_{\mathsf{v},x,y\to z})>0$. We take $\Sigma^{\prime}=\hat{\Omega}_{\mathsf{v},x,y\to z}$. It is invariant under $G_{\mathsf{v},x,z}$ and $\mathsf{H}^{\prime}=H_{\mathsf{v},x,y}$, and the former group is equal to $\mathsf{G}=G_{\mathsf{v},x}$ by Fact 2. Now take $h\in\mathsf{H}\setminus\mathsf{H}^{\prime}$. Then $h(x)=x$ and $h(y)\neq y$. Then $h\Sigma^{\prime}=\hat{\Omega}_{\mathsf{v},x,h(y)\to z}$, which is disjoint from $\hat{\Omega}_{\mathsf{v},x,y\to z}$, as desired. ∎
@@ 8.3 From commensurated to normal
Let $\mathsf{v}\in V\Gamma^{e}$. Let $\ell_{1},\ell_{2}$ be two $\mathsf{v}$-lines in the same $H_{\mathsf{v}}$-orbit. Let $A$ be any finite-index infinite cyclic subgroup of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\cap\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$.
Then $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1}):A]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2}):A]$.
@@
Recall that every $\mathsf{v}$-line $\ell$ determines a rank $1$ vertex $v_{\ell}$ in $\mathbb{B}$, and vertices of $\ell$ correspond to rank 0 vertices adjacent to $v_{\ell}$. Therefore, for every $\mathsf{v}$-line $\ell$, the action of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ on $\ell$ has finite stabilizers (Lemma 6.8) and finitely many orbits (Lemma 8.1). As $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})$ and $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$ are commensurable (Proposition 8.5 applied to $Z_{2}=\mathcal{L}_{\mathsf{v}}$), and $A$ is torsion-free and has finite index in $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\cap\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$, it follows that the actions $A\curvearrowright\ell_{1}$ and $A\curvearrowright\ell_{2}$ are free with finitely many orbits.
As $A\subseteq H_{\mathsf{v}}$, the $A$-action on $P_{\mathsf{v}}=Z_{\mathsf{v}}\times\mathcal{L}_{\mathsf{v}}$ preserves the product structure. Therefore the two actions $A\curvearrowright\ell_{1}$ and $A\curvearrowright\ell_{2}$ have the same number of orbits. Let $h\in H_{\mathsf{v}}$ be such that $h(\ell_{1})=\ell_{2}$. Then $h$ induces a bijection $h_{\ast}:\mathcal{O}_{1}^{j}\mapsto\mathcal{O}_{2}^{j}$ between the set of orbits $\{\mathcal{O}^{1}_{1},\dots,\mathcal{O}^{k}_{1}\}$ of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})\curvearrowright\ell_{1}$, and the set of orbits $\{\mathcal{O}^{1}_{2},\dots,\mathcal{O}^{k}_{2}\}$ of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})\curvearrowright\ell_{2}$. In addition $h_{\ast}$ preserves the (finite) cardinalities of stabilizers, i.e. the $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1})$-stabilizer of any point in $\mathcal{O}_{1}^{j}$ has the same cardinality (denoted by $k_{j}$) as the $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2})$-stabilizer of any point in $\mathcal{O}_{2}^{j}$. Now, for every $i\in\{1,2\}$, the number of $A$-orbits on $\ell_{i}$ is equal to $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{i}):A]\sum_{j=1}^{k}\frac{1}{k_{j}}$. Therefore $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{1}):A]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell_{2}):A]$. ∎